# EraseBench: Understanding The Ripple Effects of Concept Erasure Techniques

### v1

@@ 4.4 Experimental Setup
In our experiments, we use Stable Diffusion (SD) as the text-to-image model. We applied concept erasure techniques to each primary concept in the EraseBench dataset. For each concept, we erase a single target using both the default and reported best settings from the respective baseline papers. After erasing the target concept, we evaluate model behavior across related non-target concepts within the same dimension in EraseBench, assessing the effects of erasure on visually or semantically similar concepts. Each concept in EraseBench is represented by over 10 unique prompts, which vary in length from short, simple phrases to more complex, descriptive ones. In order to capture the diverse outputs. We generate 10 distinct image samples per prompt by varying the random seed (total of 10 random seeds). This approach provides a comprehensive view of the erasure impact across different prompts and generations, enabling us to evaluate both the intended erasure and potential ripple effects on related concepts within the same dimension.
@@
Figure 4: Erasure affects fine-grained alignment. Image generated for the prompt “A tiger perched on a rocky outcrop surrounded by mountains and a serene blue sky.” before (left) and after (right) erasing the concept cat using MACE [23]). To measure the fine-grained alignment of concepts in complex prompts, we use the visual question answering framework Gecko with question-answers generated via Gemini 1.5.
4.5 Evaluation Metrics
We use three different automated evaluation metrics to quantify the post erasure performance: CLIP [29], Rich Automatic Human Feedback (RAHF [21]), and Gecko [39]. We leverage CLIP [29] as a zero-shot classifier to determine whether a model generates the desired concept. RAHF [21] provides aesthetic and artifact score, with higher numbers denoting better aesthetics and lesser artifacts. RAHF can also provide region proposals containing artifacts in an image. Finally, we use Gecko [39] as a visual question answering based evaluation framework, with a Gemini 1.5 model to both generate the questions in the first step, and answer them for given images in the second step.
5 Results and Discussion
5.1 Concept Erasure Reduces T2I Alignment
Original Image Refer to caption
Er: Cat; Pr: Cheetah
Artifact Scores:
Original 0.98
UCE 0.78
MACE 0.83
UCE
MACE
@@ 5.1 Concept Erasure Reduces T2I Alignment
Table 1 presents the zero-shot predictions from CLIP as percentages (%). Following the definitions in [23], we measure two key attributes: Efficacy (Eff.) and Generality (Gen.). Efficacy represents the accuracy of concept erasure (accuracy on the erased class), where lower values indicate better erasure. Generality reflects accuracy on paraphrased or synonymous non-erased concepts, showing how well general information is erased. Additionally, we introduce Sensitivity (Sens.), defined as the accuracy for non-target but similar concepts. Most of the concept erasure techniques decreased accuracy on the target concept class (a desired outcome). Receler and AdvUnlearn achieved the best efficacy in erasure. Only a few techniques showed strong generality, indicating that erasure was not applied in a broad, global manner. In particular, Receler consistently displayed the best generality, preserving accuracy across target erased classes and paraphrased non-erased concepts. This may be attributed to the fact that Receler was trained using adversarial prompting and employs “erasers” designed to capture textual semantic concepts effectively. Overall, we observed a decrease in sensitivity (a non-desired outcome). This exhibits a text-to-image misalignment for non-target-non-erased concepts. There is noticeable misalignment between the text prompt and the generated image post erasure of the target concept. Low sensitivity suggests that generated images for similar, non-target concepts frequently fall out of distribution—indicating that the model has, in effect, ”learned” to overlook related concepts indirectly rather than through direct erasure. This results in images that fail to align accurately with the intended text, underscoring the model’s inability to fully retain non-target concepts after erasure is applied. In Figure 3 (row 1), we observe that when models undergo erasure of the “mouse” concept, prompting them with related concepts like ”hamster” reveals unintended effect. These erased models appear to have also ”forgotten” how to generate a hamster accurately (at least 4 out 5 techniques). This demonstrates a significant vulnerability in these erasure techniques, highlighting their limitations and the risks associated with deploying them in real-world, unrestricted settings. This trend of low sensitivity is consistent across dimensions of EraseBench. If we want evaluate overall, UCE has presented better sensitivity. Figure 4 shows an example of question answer pairs for a pair of an original and a sanitized image using Gecko [39]. While some aspects of the prompt are still correctly depicted in the image, the score is lower for sanitized model as it is missing the animal.
@@
Table 2: Text-to-image alignment scores using Gecko metric. The average scores and the standard error of the mean between prompts containing the erased and non-erased concepts and images generated with different sanitized models. Drop in score compared to the non-sanitized model shown in brackets.
@@
UCE [8] $57.6\pm 1.7\,({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-27.9})$ $74.3\pm 0.8\,({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-4.9})$
MACE [23] $38.2\pm 1.3\,({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-47.4})$ $67.9\pm 0.9\,({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-11.3})$
AdvUnlearn [47] $43.1\pm 1.4\,({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-28.0})$ $68.6\pm 0.9\,({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-10.6})$
[…]
@@
Visual Similarity (Object) Binomial Subset of superset
@@
UCE [8] 72.44 $\pm$ 0.9 73.77 $\pm$ 0.2 75.0 $\pm$ 1.0 74.19 $\pm$ 0.3 78.88 $\pm$ 3.4 78.5 $\pm$ 1.0 74.55 $\pm$ 0.6 72.20 $\pm$ 0.4
MACE [23] 73.5 $\pm$ 0.8 71.8 $\pm$ 0.07 72.0 $\pm$ 1.1 73.4 $\pm$ 1.2 72.79 $\pm$4.0 77.22 $\pm$ 0.7 72.64 $\pm$ 0.3 75.33 $\pm$ 0.2
AdvUnlearn [47] 74.6 $\pm$ 0.9 72.3 $\pm$ 1.2 69.78 $\pm$ 0.6 70.10 $\pm$ 0.9 66.23 $\pm$ 5.3 76.79 $\pm$ 1.1 71.87 $\pm$ 0.5 72.49 $\pm$ 0.2
5.2 Concept Erasure Reduces Quality For Retained Concepts.
Another important question emerges: what if the erased model retains similar non-target concepts? would this suggest that the model is unaffected by ripple effects? To explore this further, we examined the various dimensions of EraseBench. We specifically employed metrics that better align with human feedback, such as artifact measurements that quantify the degree of perceptible distortions and aesthetic quality—factors that reflect how humans perceive image fidelity. In Table 3, we present the RAHF alignment scores of individual concepts under different dimensions of EraseBench. Our results reveal a significant decline in overall quality compared to the original SD, suggesting that even when non-target concepts are preserved, the erasure process still leads to noticeable degradation in image quality. This indicates that similar and related concepts remain vulnerable to generation flaws, despite attempts at concept retention. In Figure 3, we observe that most models capable of retaining and generating the intended concept after erasure still exhibit various types of distortions. These include misaligned body parts for animal classes, cropped concepts, nonsensical text distortions, decreased size of the generated concept, and a general lack of sharpness. Additional examples can be found in the Supplemental Material. What about concepts related to art? We observe a significant drop in text-to-image alignment, highlighting a major challenge in artistic style erasure as highlighted in Table 1. Similar styles or concepts across artists are at risk of being inadvertently erased if one artist’s style is closely related to another. We observe this issue in the context of text-to-image alignment and found a marked degradation in quality, particularly in the style of non-target concepts. In Figure 3, we show the erasure of “Bosch” (the artist) and its impact on the ability to generate images of other artists, such as Altdorfer. Notably, we see a significant shift in style even for the non-erased concept. For this reason, we specifically chose not to use the RAHF alignment scores for artistic style, as the models were not trained to evaluate artistic style in line with their documentation.
@@ 5.3 Empirical Validation with Human Preferences
Figure 6: Human image preferences between images generated by the original and the erased model. The erased model used is UCE. Results show that humans prefer SD over UCE.
@@
We conduct a human preference study to additionally validate results and we obtained with automated evaluation techniques. We recruited 11 participants from our institution to judge images based on three different criteria: overall image quality, distortions, and text-to-image alignment. All participants were instructed on how to complete the task, and have provided consent to participate in the study. Each participant was presented with 50 side-by-side image comparisons, corresponding to 50 pairs of images generated by the original model and by the model with a concept erased. The order of images was randomized for each pair during presentation. The participants were asked the following three questions: ($i$) Which image exhibits superior overall quality? ($ii$) Which of the following images displays LESS noticeable distortions? ($iii$) Which image most accurately reflects and is aligned with the text label? Participants selected one of the three possible answers: Image A, Image B, Neutral. In total, we collected 1650 responses. We focus on images of non-erased concepts, as the primary hypothesis we wanted to test was whether non-target visually similar concepts were negatively affected with the erasure technique. We used UCE as it was overall the best performing technique on Sensitivity in Table 1 and also the model with the highest alignment score for non-erased concepts in Table 2. The results in Figure 6 show that humans judge images generated with the original model as having better overall quality and having fewer distortions. The original model is also preferred for better alignment, although it is closer to the “Neutral”, meaning that either both the original and sanitized images were of equally good or equally poor quality. Overall, we find that human preferences corroborate findings we observe with automatic metrics.
6 Conclusion
This work presents a thorough evaluation of concept erasure techniques, uncovering shortcomings in their robustness and reliability. We introduce EraseBENCH, a comprehensive benchmark that reveals critical gaps in current methods, particularly when faced with visually similar, binomial, and semantically related concepts. Our findings demonstrate that despite recent advancements, existing techniques struggle to preserve model quality and ensure consistent erasure across a variety of scenarios. This highlights the need for more robust and nuanced approaches to concept erasure, as well as more rigorous evaluation protocols and metrics.

### v2

@@ 4 Experimental setup
Experiments. We use Stable Diffusion (SD) as our T2I model and apply concept erasure techniques to each primary concept in EraseBench. For each concept, we perform erasure using both the default and best-reported settings from baseline methods. We maintain the default retain sets (e.g. COCO prompts for AdvUnlearn) and, for methods requiring anchor concepts, we map them to an empty concept (”). We evaluate the sanitized models on related non-target concepts within the same EraseBench dimension, assessing the effects on visually and semantically similar concepts. Each concept is represented by over 10 unique prompts, ranging from simple to complex descriptions. To capture diverse outputs, we generate 10 images per prompt with different random seeds. This approach provides a comprehensive evaluation of erasure effectiveness and its unintended effects on related concepts.
Evaluation Metrics. We use different types of automated evaluation metrics to quantify and explain the detrimental effect erasure techniques can have on the quality of generated images of non-target concepts. We use three different automated evaluation metrics to quantify the post erasure performance: CLIP [36], Rich Automatic Human Feedback (RAHF [28]), and Gecko [48]. CLIP is a widely used joint image and text embedding model which quantifies the similarity between text and image representations. We leverage CLIP as a zero-shot classifier to determine whether a model generates the desired concept. The RAHF model is trained on rich human feedback for generated images with a range of scores to evaluate overall image quality, and also mark points about which regions are problematic. The RAHF quality scores include an aesthetic score, which assesses the visual appeal of an image, and an artifact score (i.e. plausibility score), which gauges the presence and extent of visual artifacts. We use both the aesthetic and artifact scores as measurements of quality for generated images, following [28], where higher scores indicate better quality (i.e. good aesthetics and minimal artifacts). In addition to scores, the RAHF model provides heatmaps for each image to highlight artifact-prone areas, which we use for visual inspection in our benchmark. Finally, we use a Visual Question Answering (VQA)-based metric called Gecko, a VQA evaluation framework. In our setup, we leverage the multi-modal Gemini 1.5 model for both steps: first, to generate text-related questions, and second, to answer them based on the given image, similar to other VQA methods [2, 51]. One advantage of such metrics is that they do not require task-specific fine-tuning and provide an interpretable alignment score. This evaluation can be replicated using other available models. Typically, an LLM generates questions related to the text prompt in the first step, while a vision-language model (VLM) answers those questions based on the generated image. The final score is computed as the average of correctly answered questions, allowing for backtracing to identify which aspects of the text are misaligned with the generated image.
5 Results
5.1 Concept Erasure Reduces T2I Alignment For Non-Target Concepts (Over-Erasure)
@@
Figure 4: Erasure affects fine-grained alignment. Prompt: “A tiger perched on a rocky outcrop surrounded by mountains and a serene blue sky.” before (left) and after (right) erasing the concept cat using MACE [30]).
@@
Table 1 presents the zero-shot predictions from CLIP as percentages (%). Following the definitions in [30], we measure two key attributes: Efficacy (Eff.) and Generality (Gen.). Efficacy represents the accuracy of concept erasure (accuracy on the erased class), where lower values indicate better erasure. Generality reflects accuracy on paraphrased or synonymous non-erased concepts, showing how well general information is erased. Additionally, we introduce Sensitivity (Sens.), defined as the accuracy for non-target but similar concepts. Most of the concept erasure techniques decreased accuracy on the target concept class (a desired outcome). Receler, MACE, and AdvUnlearn achieved the best efficacy in erasure overall per EraseBench dimension. Only a few techniques showed strong generality, indicating that erasure was not applied in a broad, global manner. In particular, Receler consistently displayed the best generality, preserving accuracy across target erased classes and paraphrased non-erased concepts. This may be attributed to the fact that Receler was trained using adversarial prompting and employs “erasers” designed to capture textual semantic concepts effectively. Overall, we observed a decrease in sensitivity, an undesirable outcome that highlights T2I misalignment for non-target non-erased concepts, essentially, over-erasure. After erasing a target concept, we noticed a clear mismatch between the text prompt and the generated image for related non-target concepts. Low sensitivity suggests that instead of selectively erasing only the target concept, the model inadvertently suppresses related concepts, causing generated images to fall out of distribution. This indicates that the model has not only erased the intended concept but also learned to overlook semantically or visually similar ones. As a result, the generated images fail to align with the intended prompts, revealing the model’s inability to fully preserve non-target concepts after erasure. In Figure 2 (row 1), we observe that when models undergo erasure of the koala concept, prompting them with related concepts like Tree Kangaroo reveals unintended effect. These erased models appear to have also forgotten how to generate a Tree Kangaroo accurately. Similarly, when we erase the artist concept Bosch and prompt the sanitized model to generate artwork in the style of Altdorfer, the model struggles to capture the distinctive artistic style of the non-erased concept. This shows that the erasure process affects not only the target concept but also related artistic styles, leading to a loss of fidelity in the generated images. This demonstrates a significant vulnerability in these erasure techniques and highlights their risks associated with deploying them in real-world unrestricted settings. This trend of low sensitivity is consistent across dimensions of EraseBench. If we want to evaluate overall, UCE demonstrates superior preservation of visual concept components compared to other methods.
Table 2: Harmonic mean ($\uparrow$) of efficacy, generality, and sensitivity scores on explicit concepts across EraseBench dimensions. Each column represents a distinct relationship type: visually similar objects, artist style similarity, subset-superset relationships, and binomial pairs. Higher scores indicate better balance between erasure success and minimal unintended degradation. While all methods outperform the original model, performance varies by dimension, revealing trade-offs in handling sensitive and safety-critical content.
Technique Visual Sim. Subset-superset Binomial
Original 16.37 23.48 19.64
UCE [11] 63.16 62.62 76.13
AdvUnlearn [54] 67.66 58.28 71.33
Receler [18] 68.47 56.31 74.88
@@
Table 3: Text-to-image alignment scores using the Gecko metric. The average scores and the standard errors of the mean for images generated with different techniques and prompts that either contain the erased concept or not. ”Original” is the baseline model w/o any concept erasure. Drop in score compared to the original model shown in brackets.
@@
UCE [11] $57.6\pm 1.7\,({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-26.4})$ $74.3\pm 0.8\,({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-3.4})$
MACE [30] $38.2\pm 1.3\,({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-45.9})$ $67.9\pm 0.9\,({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-9.8})$
AdvUnlearn [54] $43.1\pm 1.4\,({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-41.0})$ $68.6\pm 0.9\,({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-9.0})$
Original Image Refer to caption
Er: Goldfish
Pr: A Guppy
Artifact Scores:
Original:93.63
UCE: 84.21
AdvUnlearn:79.01
UCE
AdvUnlearn
Figure 5: Erasure introduces artifacts during subset-superset concept generation. We erase concept ”goldfish” and generate images for the prompt ”an image of a guppy”. We present the RAHF artifact heatmaps for images generated post-erasure via AdvUnlearn and UCE. We see that the artifact introduced by each method can vary spatially and by intensity, which prompts our inclusion of the artifact score in EraseBench. More heatmaps examples can be found in the supplemental material.
Table 2 presents the harmonic mean of efficacy, generality, and sensitivity scores for explicit concepts (see Table 8 in the supplemental material for details) across three EraseBench dimensions: Visual Similarity, Subset-superset, and Binomial. UCE demonstrates consistently strong performance across most dimensions, while AdvUnlearn and Receler achieve particularly high scores in Subset-superset and visual similarity settings. These results highlight the value of evaluating concept erasure across diverse relational categories to ensure method robustness and general applicability.
Next, we use Gecko to evaluate image-text alignment with a VQA-based approach. Unlike CLIP, Gecko provides an interpretable alignment score as it is computed over question-answer pairs. Figure 4 shows an example of question-answer pairs for a pair of an original and a sanitized image. While some aspects of the prompt are still correctly depicted in the image, the score is lower for sanitized model as it is missing the animal. We compute alignment scores between images generated by original and sanitized models (UCE, MACE, and AdvUnlearn) and prompts corresponding to four different concepts (goat, goldfish, cat, and Vincent Van Gogh) from four different dimensions. In total, we compute alignment scores for 6246 text-image pairs. The results in Table 3 show an expected drop in scores for images generated with sanitized models compared to images generated with the original model (c.f. column “Erased concepts”), as we expect the erased concept to be missing in images generated by sanitized models. However, we also observe a drop in scores for non-erased concepts across different dimensions. While that drop is much smaller, it is statistically significant across all three comparisons ($\alpha=.01$ using the Wilcoxon signed-rank test). The UCE technique has the smallest gap with the original model for non-erased concepts, while MACE has the largest. The results for UCE are consistent with CLIP-score zero-shot prediction in Table 1, where UCE was consistently the technique with highest sensitivity across all dimensions. We also manually inspected a subset of questions, answers and images, to ensure that the lower score is not caused by the VQA model latching on spurious correlations, but is indeed due to depicted concepts missing or the image being of lower quality. We observe that while some concepts are generally more difficult to depict correctly even in the original model (e.g. goldfish), this is reflected in the lower overall alignment score for the original model (e.g., for all three techniques the alignment score of the original model is in 60s for the concept goldfish, while it is in 80s and 90s for other concepts). Finally, question-answer breakdown in Gecko allows us to pinpoint some failures of image generation in sanitized models. When looking at answers to questions asking about object presence (eg., Is there an $\langle$ object $\rangle$ in the image?), we found that for all three techniques erased models had lower Gecko scores compared to original models on prompts with visually similar objects (UCE: 80.3% vs 89.0%, MACE: 67.1% vs 89.0%, and AdvUnlearn: 83.1% vs 89.0%), which highlights unintentional consequences of concept erasure techniques.
5.2 Concept Erasure Reduces Quality For Non- Target Concepts
@@
Visual Similarity (Object) Binomial Subset- superset
@@
UCE [11] 72.44 $\pm$ 0.9 73.77 $\pm$ 0.2 75.00 $\pm$ 1.0 74.19 $\pm$ 0.3 78.88 $\pm$ 3.4 78.50 $\pm$ 1.0 74.55 $\pm$ 0.6 72.20 $\pm$ 0.4
MACE [30] 73.50 $\pm$ 0.8 71.80 $\pm$ 0.07 72.00 $\pm$ 1.1 73.40 $\pm$ 1.2 72.79 $\pm$4.0 77.22 $\pm$ 0.7 72.64 $\pm$ 0.3 75.33 $\pm$ 0.2
AdvUnlearn [54] 74.60 $\pm$ 0.9 72.30 $\pm$ 1.2 69.78 $\pm$ 0.6 70.10 $\pm$ 0.9 66.23 $\pm$ 5.3 76.79 $\pm$ 1.1 71.87 $\pm$ 0.5 72.49 $\pm$ 0.2
Global Concepts. Beyond non-target misalignment, T2I models also exhibit other forms of degradation post-erasure, such as distortions (i.e. artifacts) and poor aesthetics. To explore this further, we specifically employed metrics that better align with human feedback, such as artifact measurements that quantify the degree of perceptible distortions and aesthetic quality (i.e. factors that reflect how humans perceive image fidelity). In Table 4, we present the RAHF alignment scores of individual concepts under different dimensions of EraseBench. Our results reveal a significant decline in overall quality compared to the original SD, suggesting that the erasure process leads to noticeable degradation in image quality. This indicates that similar and related concepts remain vulnerable to generation flaws, despite attempts at concept retention. In Figure 2, we observe that most models capable of retaining and generating the intended concept after erasure still exhibit various types of distortions. These include misaligned body parts for animal classes, cropped concepts, nonsensical text distortions, decreased size of the generated concept, and a general lack of sharpness. Additional examples can be found in the Supplemental Material. Artists Concepts. We observe a significant drop in T2I alignment, highlighting a major challenge in artistic style erasure as highlighted in Table 1. Similar styles or concepts across artists are at risk of being inadvertently erased if one artist’s style is closely related to another. We observe this issue in the context of T2I alignment and found a marked degradation in quality, particularly in the style of non-target concepts. In Figure 2, we show the erasure of Degas (the artist) and its impact on the ability to generate images of other artists, such as Cassatt. Notably, we see a significant shift in style even for the non-erased concept. We specifically chose not to use the RAHF alignment scores for artistic style, as the models were not trained to evaluate artistic style based on their documentation.
@@ 5.3 Empirical Validation with Human Preferences
We conduct a human preference study to additionally validate results obtained with automated evaluation techniques. We recruited 11 participants from our institution to judge images based on three different criteria: overall image quality, distortions and T2I alignment. All participants were instructed on how to complete the task, and have provided consent to participate in the study. Each participant was presented with 50 side-by-side image comparisons, corresponding to 50 pairs of images generated by the original model and by the model with a concept erased. The order of images was randomized for each pair during presentation. The participants were asked the following three questions: ($i$) Which image exhibits superior overall quality? ($ii$) Which of the following images displays LESS noticeable distortions? (iiiiii) Which image most accurately reflects and is aligned with the text label?
Participants selected one of the three possible answers: Image A, Image B, or Neutral. In total, we collected 1650 responses for UCE and 1485 for AdvUnlearn. We focus on images of non-erased concepts, as the primary hypothesis we wanted to test was whether non-target visually similar concepts were negatively affected with the erasure technique. The results in Figure 6 show that humans judge images generated with the original model as having better overall quality and having fewer distortions. The original model is also preferred for better alignment, although it is closer to the “Neutral” for UCE, meaning that either both the original image and sanitized image were of equally good or equally poor quality. We provide additional human preferences for AdvUnlearn in the supplemental material. Results in Figure 10 also show that most participants preferred the original images with respect to quality, alignment, and less artifacts. Overall, we find that human preferences corroborate findings we observe with automatic metrics.
Figure 6: Human image preferences between images generated by the original and the erased model. The erased model used is UCE. Results show that humans prefer SD over UCE. We provide the results of AdvUnlearn in the supplemental material.
6 Discussion
In our work, we have identified and emphasized critical issues surrounding T2I models post concept erasure, making it our primary focus. Addressing these challenges is the logical next step, but it raises a complex research question that demands further exploration. In this section, we discuss into possible mitigation strategies and assess their impacts. We stress that finding effective solutions is an ongoing, non-trivial process that requires continued investigation and innovation.
The Role of Retain Sets in Mitigating Post-Erasure Distortions. A key question is whether the design of the retain set, specifically including concepts closely intertwined with the target erased concept can help alleviate post-generation distortions. We try to investigate whether adding non-target concepts that share semantic or visual similarities can reduce unintended side effects caused by erasure. To explore this, we performed erasure on targeted concepts using EraseBench while strategically incorporating related non-target concepts into the retain set. For example, when erasing cat, we progressively introduce related concepts such as tiger and panther into the retain set and evaluate the impact on both generated outputs and artifact persistence. We show in Figure 7 the influence of the retain set on non-target concepts. We compare the original SD without erasure, erased model without a retain set (i.e. no explicit retain set), and erased model with a retain set in which we introduce the entangled non-target concepts one by one.
(i) Findings on non-target concepts. Our investigation indicates that incorporating entangled non-target concepts into the retain set partially alleviates over-erasure (misalignment on non-target concepts), preserves aspects of related non-target concepts. However, artifacts and distortions persists, particularly in non-targeted non-erased concepts suggesting that while carefully designing a retain set may help maintain concept structure, it does not fully resolve the unintended distortions.
(ii) Findings on target-erased concepts. Examining the erased target concepts in Figure 8, we find that incorporating multiple entangled concepts into the retain set helps alleviate some post-erasure artifacts on the non-target concepts but comes at the cost of concept leakage. Specifically, this leakage manifests as spurious regeneration, where the erased concept partially or fully re-emerges in generated outputs. Our observations suggest that while including related non-target concepts in the retain set can mitigate over-erasure, it increases the risk of unintended regeneration. The more entangled concepts added, the greater the likelihood of leakage, highlighting a trade-off between preserving non-target concepts and ensuring effective erasure.
The Impact of Anchor Sets on Mitigating Unintended Effects of Concept Erasure. We investigate whether anchor concepts can help in mitigating the unintended distortions in non-target concepts observed after concept erasure. An anchor may act as a reference point, defining boundaries for selective erasure while preserving related concepts. To test this, we conducted experiments in Figure 9 where we erased Van Gogh while introducing Post-Impressionism as an anchor concept. Similarly, we erased Grumpy Cat while anchoring the model with Internet meme, as outlined in [8]. The intuition behind this approach is that a well-defined anchor might help the model retain a coherent representation of related concepts, thereby reducing unintended consequences of erasure. We observed models struggle with generating good quality non-target concept images, suggesting anchors do not consistently improve stability. While anchors may help in alignment, they do not fully resolve artifacts or style preservation issues, warranting further investigation into their effectiveness.
Multi-Concept Erasure: Intra-type and Inter-type Effects. We also investigate multi-concept erasure, differentiating between intra-type (main and paraphrased concepts) and inter-type (main and unrelated concepts) erasure. Average artifact scores in the visual similarity domain reach 78.3% with intra-type and 71.4% with inter-type erasure. The former outperforms the single concept erasure at 71.9% (see example in Figure 15 in the supplemental material). Our findings show that intra-type erasure (i.e. removing multiple related concepts) more effectively eliminates targets and reduces artifacts when faced with non-target but related prompts. In contrast, inter-type erasure offers minimal artifact reduction, suggesting that carefully grouping related concepts enhances erasure precision without degrading generation quality.
Figure 7: Impact of adding visually similar concepts to the retain set with respect to non-target concepts. Without a retain set, image quality degrades significantly. Unrelated concepts (e.g. house, elephant, and boat) fail to prevent artifacts, while a curated, visually similar retain set improves quality but remains below pre-erasure levels.
Figure 8: Impact of adding visually similar concepts to the retain set with respect to the target concept. Without a retain set, erasure is effective, but as similar concepts are added, the erased concept gradually re-emerges.
Figure 9: Impact of Anchor Sets on Mitigating Unintended Effects of Concept Erasure.
7 Conclusion and Limitations
In this work, we present a thorough evaluation of concept erasure techniques and uncover shortcomings in their reliability. By introducing EraseBench, we reveal critical gaps in current methods. Our findings suggest that despite recent advancements, existing concept erasure techniques fall short in preserving model quality, which call for the need for more robust and nuanced approaches as well as more robust evaluation protocol and metrics. While our framework provides a diverse evaluation, automating concept selection could improve scalability. Given dimensions like binomiality, researchers can explore broader concept spaces beyond our curation. Additionally, a deeper exploration of optimal retain or erasure sets to minimize distortions remains an open question. We hope EraseBench inspires new research directions toward developing more reliable concept erasure techniques.